# XUE: The CO2-rich terrestrial planet-forming region of an externally irradiated Herbig disk

### v1

@@ 5 Discussion
5.1 XUE 10’s peculiar spectrum
Among all the MIRI MRS spectra of protoplanetary disks published so far (see for e.g. the MINDS and JDISC programs: Henning et al., 2024; Arulanantham et al., 2025), XUE 10 is the first Herbig disk located in a massive star-forming region to be analyzed. The comparison with MINDS and JDISC published sources, including the ones discussed in Sec. 3 (i.e., GW Lup, CX Tau, Sz 98, MY Lup) must be taken with caution, since they all are low-mass to solar-mass T Tauri disks located in low-mass isolated star-forming regions. Besides the striking simultaneous detection of four CO2 isotopic species in emission, the mid-IR spectrum of XUE 10 appears to be ?line-poor?, i.e, with non-detections of H2O, OH, C2H2, HCN etc., in contrast to the typical ?line-rich? mid-IR spectrum of T Tauri sources. The ?line-poor? aspect of XUE 10 is in common with other known IMTT and Herbig disk sources. The origin of this discrepancy between T Tauri and disks surrounding intermediate-mass stars remains unknown. However, we do have samples of Herbig disks in isolation and in massive star-forming regions (e.g., Stapper et al., 2022, 2025), but we lack the JWST/MIRI spectra to do a proper comparison, and establish if there are systematic differences in the gas content of these two groups of disks. Moreover, the brightness of IMTT/Herbig young stars, which tend to saturate in nearby regions, and the consequent higher disk infrared continuum, have often limited the observational analysis of the inner disk regions to the discussion of dust features rather than of the gas-phase molecular chemistry.
The main references for the latter are studies of nearby ( 700 pc) Herbig disks observed at 1–5 m with VLT/CRIRES (Fedele et al., 2011), 4.5–5.2 m with the iSHELL instrument at the Infrared Telescope Facility (Banzatti et al., 2022), and with Spitzer at 10–36 m (Banzatti et al., 2020). The general finding in these samples is the lack of the chemical richness that is seen in the atmosphere of T Tauri disks, with mostly upper limits or detections with line fluxes 10-14 erg cm-2 s-1 of H2O, OH, 12CO, 13CO, CO2, HCN and C2H2. Interestingly, at the longer wavelengths probed by Herschel (50–200 m), H2O and OH are detected with similarly low column densities of 1014–1015 cm-2 at large emitting radii ( 10 au), pointing at an outer–inner depletion of water molecules in Herbig disks by means of the UV irradiation from the central massive protostars (Fedele et al., 2011; Heays et al., 2017). XUE 10 therefore looks similar to other IMTT/Herbig disks in the sense of a bright dust continuum with PAHs and silicate dust emission, and because of the lack of a ?variety? of bright gas-phase molecular species.
@@ 5.1 XUE 10’s peculiar spectrum
In Figure 7 we compare the line luminosities of H2O, CO2, CO and the 13–25 m spectral index observed in XUE 10 with the rest of the XUE sample (Ramírez-Tannus et al., 2025), the JWST observations of the CO2-rich GW Lup T Tauri disk in the nearby Lupus I cloud ( 155 pc; Grant et al., 2023), and with Spitzer measurements of other T Tauri and Herbig PPDs located in nearby ( 200 pc), low-mass star-forming regions ( 3 Myr), adopting the samples from Banzatti et al. (2020, 2022). The line luminosities are estimated from the observed spectra at 17.08–17.4 m (H2O), where for XUE 10 we derive an upper limit in the CO2 residual spectrum, 14–15 m (12CO2), and 4.9–5.2 m (CO). We find that XUE 10 stands out for the highest 12CO2/H2O line luminosity ratio, which must be interpreted as a lower limit. For CO, XUE 10 is at the high end of the distribution, but not such a strong outlier, especially in comparison to the other XUE sources. Following the approach of Ramírez-Tannus et al. (2023), which was originally presented in Perotti et al. (2023), we also computed the continuum spectral index between 13 and 25 m, as a proxy of the level of dust depletion of the inner disk. The rightmost panel of Figure 7 suggests that the inner disk of XUE 10 is not depleted in m-size dust grains, similarly to the bulk of surveyed T Tauri and Herbig disks, albeit a large spread is present. Dust depletion of pebble-size grains is instead common among disks observed around Herbig stars (e.g., Banzatti et al., 2018), but with the present data this cannot be assessed for the XUE sample.
Given the relatively strong external radiation field incident on XUE 10, it is also worth comparing our results with the d203-506 proplyd in the ONC. Notably, we do not detect the ro-vibrational band of CH3+ at 7.15 m, which has first been detected in the photo-evaporating wind of d203-506 (Berné et al., 2023), and later in the non-irradiated TW Hya disk (Henning et al., 2024). This non-detection may indicate that XUE 10 lacks a photo-dissociation region (PDR), which has been proposed as the origin of CH3+ emission observed towards d203-506, i.e., in the extended PDR surrounding the disk, as well as in the irradiated edge of the Orion Bar (Zannese et al., 2025). This may be explained by the difference in FUV exposure between XUE 10 ( 103 G0) and d203-506 ( 104 G0), or shielding of XUE 10 from external irradiation, or the possibility that external photoevaporation has already removed most of its outer disk (Ramírez-Tannus et al., 2025). In the non-irradiated case of TW Hya, instead, the photo-dissociation region producing CH3+ is interpreted as a gas depleted cavity, creating the PDR-like emission region. Additionally, OH emission, which has been proposed to signpost the destruction of water by means of FUV radiation in d203-506 (Zannese et al., 2024), is also absent in XUE 10’s spectrum. A more detailed discussion of the OH non-detection in XUE 10 is provided in the following sections.
@@ 5.2 Inner disk chemistry scenarios
We can summarize the presence and replenishment of CO2 and H2O in the inner disk of protoplanetary disks via two mechanisms: ice sublimation of the mantle from radially drifting dust grain at the H2O and CO2 ice lines, and local gas-phase formation through hydroxide (OH). In particular, interstellar and cometary ices are known to be rich in CO2, with measured abundances of 15–25% compared to H2O ice, or 10-5 with respect to the total gas density (e.g., de Graauw et al., 1996; Bergin et al., 2005; Boogert et al., 2015; Le Roy et al., 2015). We neglect grain-surface formation of CO2 in the inner disk, which is negligible compared to gas-phase production (Bosman et al., 2017). Considering the latter, OH can react with CO to form CO2 via: CO + OH CO2 + H (Smith et al., 2004), which has a negligible activation energy and is most effective at gas temperatures up to 400 K, after which destruction channels for OH and CO2 start to become important. In fact, OH can also react with molecular hydrogen to form water: OH + H2 H2O + H (Glassgold et al., 2009). This reaction becomes effective at 200–300 K, due to a high activation barrier of nearly 1700 K, necessary to form OH first (McElroy et al., 2013). Given the exothermic (spontaneous) nature of the CO2 and H2O formation channels, and the availability of gas-phase H2 and OH reactants in the warm disk atmosphere (200–1000 K), H2O should be more abundant than CO2 in the inner disk, but on condition of neglecting photochemical effects and vertical/turbulent mixing. This was first evident from Spitzer observations, where a low CO2 abundance, 10-4 with respect to H2O ( 10-8 with respect to the total gas density) was found (Salyk et al., 2008; Pontoppidan & Blevins, 2014), and it has been repeatedly confirmed by JWST observations in nearby regions (e.g., Banzatti et al., 2025).
@@
There is, however, mounting evidence of CO2-rich inner disk spectra, at least among T Tauri sources (Gasman et al., 2023; Grant et al., 2023; Vlasblom et al., 2025; Salyk et al., 2025), alongside modeling efforts (Bosman et al., 2022; Sellek et al., 2025) investigating the origin of variations in the spectra of CO2 in contrast to water. XUE 10 is a special case, exhibiting the highest contrast between CO2 and H2O mid-infrared emission to date. It is currently also the only Herbig disk with a CO2-rich mid-infrared spectrum. We therefore need to explain why XUE 10 lacks bright H2O vapor emission, and how CO2 vapor can become so abundant and dominate over H2O at the disk gas temperatures and locations probed at mid-infrared wavelengths. Additionally, we aim to investigate the origin of the potential detection of oxygen isotopic anomalies in CO2 vapor in the inner disk, and contextualize our findings with the evolution of disks around young, intermediate-mass stars (⋆ 1.5 ⊙), and the external environment. We propose the following two explanatory scenarios, and discuss their application to the case of XUE 10:
@@
The molecular emission observable in the inner disk is determined by in-situ processes, i.e., stellar UV irradiation and local gas-phase chemistry.
@@
The observable inner disk emission strongly depends on radial transport processes of gas and dust from the outer disk, whose outcome is not reset by in-situ processes. These include radial drift of ice-rich dust grains and gas advection, but also processes that can impede their radial transport, specifically outer disk truncation (e.g., via external UV photoevaporation), and dust traps/gas cavities.
5.2.1 In-situ chemistry
@@
The harsher UV irradiation coming from a central Herbig star, compared to a cooler T Tauri star, can drive substantial photo-dissociation of water according to the reaction: H2O + OH + H (Heays et al., 2017). If this reaction proceeds on relatively short timescales compared to vertical/turbulent mixing in the disk and compared to the local gas-phase H2O production, water is unable to self-shield and its abundance lowers down, while the freed OH radicals can be sequestrated by CO to form CO2. CO2 itself can undergo photo-dissociation, releasing CO and O via: CO2 + CO + O (Heays et al., 2017), and O2 secondarily (Lu et al., 2014). If the UV photo-dissociation rate of H2O and CO2 molecules were of the same order of magnitude, both species would efficiently self-shield. In this sense, the CO2–H2O dichotomy could be explained by the disk thermal structure. In other words, H2O and CO2 (and OH) would get efficiently destroyed in the upper warm disk layers by stellar irradiation, while in the deeper and cooler layers atomic O and OH molecules would rather form CO2 instead of water (Bosman et al., 2018). However, both H2O and OH (with much fewer detections of CO2) have been observed towards several disks around older A- and B-type Herbig stars in nearby regions (Fedele et al., 2011; Banzatti et al., 2020), where the central stars have comparable luminosities and likely emit stronger UV radiation than the F-type star in XUE 10. We therefore speculate that CO2 photo-dissociation is not as significant as H2O photo-dissociation in XUE 10, and that CO2 but not water may self-shield in its inner regions. This implies that alternatively to thermal effects, a high column density of CO2 can be achieved in the inner disk through gas chemistry by actively removing water. Interestingly, this possibility agrees with a thermochemical DALI synthetic spectrum where water self-shielding is not included (Bosman et al., 2022), pointing towards a deep atmospheric layer of CO2. The gas temperature derived in this work for 12CO2 ( 370 K; see Table 2) together with the observation of its hot -branch bandheads at 13.9 and 16.2 m (excited only in very dense gas), further support the hypothesis of a deep CO2 layer in the disk atmosphere of XUE 10. We note that 12CO2 may be partially non-LTE excited, particularly in the -branch. Non-LTE excitation conditions may arise due to infrared pumping fields, which are known to affect other species like CO and hydrogen cyanide (HCN) at infrared wavelengths (e.g., Thi et al., 2011; Bruderer et al., 2015). This would imply a lower gas temperature and a larger emitting area than estimated under LTE assumptions. A colder gas reservoir (gas 300 K) would in turn strengthen the hypothesis of an efficient CO2 gas-phase production in the inner disk at the expense of water (Bosman et al., 2017; Salyk et al., 2025).
@@
It is important to underline that the production of CO2 vapor through its gas-phase formation channel crucially depends on the availability of OH molecules. This radical primarily forms through the reaction O + H2 OH + H, with an activation barrier of 1700 K (McElroy et al., 2013), through water photo-dissociation, and in a minor fraction via H + O2 OH + H, which is characterized by a much higher activation energy of 8000 K (Baulch et al., 1992). The non-detection of OH in XUE 10 further suggests a low water column density in its inner disk, which may limit the photo-dissociation-driven replenishment of this radical (Zannese et al., 2024). The rotationally excited emission lines of OH may also remain undetectable with JWST due to limited sensitivity above the dust continuum.
@@
This in-situ scenario can explain a low H2O column density, the non-detection of OH, and the observation of a high column density of CO2, and therefore the detection of bright CO2 isotopologues in emission. However, as water must be removed to allow efficient CO2 production in this scenario, it critically depends on the relative timescales of H2O photo-dissociation and of vertical mixing in the inner disk of XUE 10. Concerning the latter, observational evidence suggests weak turbulence in the well-studied Herbig disk HD 163296 (Flaherty et al., 2015), with no significant differences in turbulent mixing strength reported between Herbig and T Tauri disks (Mulders & Dominik, 2012). The oddity might therefore regard the timescale over which H2O photo-dissociates, which must be unusually short in XUE 10 in comparison to Herbig disks of earlier spectral type (A and B), and more evolved ( 3 Myr). Moreover, the presence of potential oxygen isotopic anomalies in CO2 vapor in the inner disk cannot be explained solely by local gas-phase channels, unless a substantial dust shielding is present in the inner disk, preventing the UV photo-dissociation of the optically thinner CO2 isotopologues forming in the gas-phase. We propose that a more realistic scenario should include the impact of radial transport processes of gas and ice-rich dust grains from the outer to the inner disk, which we consider in the following section.

### v2

@@ 5 Discussion
5.1 The peculiar spectrum of XUE 10
Among all the MIRI MRS spectra of protoplanetary disks published so far (see, e.g., the MINDS and JDISC programs: Henning et al., 2024; Arulanantham et al., 2025), XUE 10 is the first Herbig disk located in a massive star-forming region to be analyzed. The comparison with MINDS and JDISC published sources, including the ones discussed in Sec. 3 (i.e., GW Lup, CX Tau, Sz 98, MY Lup) must be taken with caution, since they all are low-mass to solar-mass T Tauri disks located in low-mass isolated star-forming regions. Besides the striking simultaneous detection of four CO2 isotopic species in emission, the mid-IR spectrum of XUE 10 appears to be ?line-poor?, i.e, with non-detections of H2O, OH, C2H2, HCN etc., in contrast to the typical ?line-rich? mid-IR spectrum of T Tauri sources. The ?line-poor? aspect of XUE 10 is in common with other known IMTT and Herbig disk sources. The origin of this discrepancy between T Tauri and disks surrounding intermediate-mass stars remains unknown. However, we do have samples of Herbig disks in isolation and in massive star-forming regions (e.g., Stapper et al., 2022, 2025), but we lack the JWST/MIRI spectra to do a proper comparison, and establish if there are systematic differences in the gas content of these two groups of disks. Moreover, the brightness of IMTT or Herbig young stars, which tend to saturate in nearby regions, and the consequent higher disk infrared continuum, have often limited the observational analysis of the inner disk regions to the discussion of dust features rather than of the gas phase molecular chemistry.
The main references for the latter are studies of nearby ( 700 pc) Herbig disks observed at 1–5 m with VLT/CRIRES (Fedele et al., 2011), 4.5–5.2 m with the iSHELL instrument at the Infrared Telescope Facility (Banzatti et al., 2022), and with Spitzer at 10–36 m (Banzatti et al., 2020). The general finding in these samples is the lack of the chemical richness that is seen in the atmosphere of T Tauri disks, with mostly upper limits or detections with line fluxes 10-14 erg cm-2 s-1 of H2O, OH, 12CO, 13CO, CO2, HCN and C2H2. Interestingly, at the longer wavelengths probed by Herschel (50–200 m), H2O and OH are detected with similarly low column densities of 1014–1015 cm-2 at large emitting radii ( 10 au), pointing at an outer–inner depletion of water molecules in Herbig disks by means of the UV irradiation from the central massive protostars (Fedele et al., 2011; Heays et al., 2017). XUE 10 therefore looks similar to other IMTT or Herbig disks in the sense of a bright dust continuum with PAH and silicate dust emission and because of the lack of a ?variety? of bright gas phase molecular species.
@@ 5.1 The peculiar spectrum of XUE 10
In Figure 7 we compare the line luminosities of H2O, CO2, CO and the 13–25 m spectral index observed in XUE 10 with the rest of the XUE sample (Ramírez-Tannus et al., 2025), the JWST observations of the CO2-rich GW Lup T Tauri disk in the nearby Lupus I cloud ( 155 pc; Grant et al., 2023), and with Spitzer measurements of other T Tauri and Herbig protoplanetary disks located in nearby ( 200 pc), low-mass star-forming regions ( 3 Myr), adopting the samples from Banzatti et al. (2020, 2022). The line luminosities are estimated from the observed spectra at 17.08–17.4 m (H2O), where for XUE 10 we derive an upper limit in the CO2 residual spectrum, 14–15 m (12CO2), and 4.9–5.2 m (CO). We find that XUE 10 stands out for the highest 12CO2/H2O line luminosity ratio, which must be interpreted as a lower limit. For CO, XUE 10 is at the high end of the distribution, but not such a strong outlier, especially in comparison to the other XUE sources. Following the approach of Ramírez-Tannus et al. (2023), which was originally presented in Perotti et al. (2023), we also computed the continuum spectral index between 13 and 25 m, as a proxy of the level of dust depletion of the inner disk. The rightmost panel of Figure 7 suggests that the inner disk of XUE 10 is not depleted in m-size dust grains, similarly to the bulk of surveyed T Tauri and Herbig disks, albeit a large spread is present. Dust depletion of pebble-size grains is instead common among disks observed around Herbig stars (e.g., Banzatti et al., 2018), but with the present data this cannot be assessed for the XUE sample.
Given the relatively strong external radiation field incident on XUE 10, it is also worth comparing our results with the d203-506 proplyd in the ONC. Notably, we do not detect the ro-vibrational band of CH3+ at 7.15 m, which has first been detected in the photo-evaporating wind of d203-506 (Berné et al., 2023), and later in the non-irradiated TW Hya disk (Henning et al., 2024). This non-detection may indicate that XUE 10 lacks a photo-dissociation region (PDR), which has been proposed as the origin of CH3+ emission observed toward d203-506, i.e., in the extended PDR surrounding the disk, as well as in the irradiated edge of the Orion Bar (Zannese et al., 2025). This may be explained by the difference in FUV exposure between XUE 10 ( 103 G0) and d203-506 ( 104 G0), or shielding of XUE 10 from external irradiation, or the possibility that external photoevaporation has already removed most of its outer disk (Ramírez-Tannus et al., 2025). In the non-irradiated case of TW Hya, instead, the photo-dissociation region producing CH3+ is interpreted as a gas depleted cavity, creating the PDR-like emission region. Additionally, OH emission, which has been proposed to signpost the destruction of water by means of FUV radiation in d203-506 (Zannese et al., 2024), is also absent in the spectrum of XUE 10. A more detailed discussion of the OH non-detection is provided in the following sections.
@@ 5.2 Inner disk chemistry scenarios
We can summarize the presence and replenishment of CO2 and H2O in the inner disk of protoplanetary disks via two mechanisms: ice sublimation of the mantle from radially drifting dust grain at the H2O and CO2 ice lines and local gas phase formation through hydroxide (OH). In particular, interstellar and cometary ices are known to be rich in CO2, with measured abundances of 15–25% compared to H2O ice, or 10-5 with respect to the total gas density (e.g., de Graauw et al., 1996; Bergin et al., 2005; Boogert et al., 2015; Le Roy et al., 2015). We neglect grain-surface formation of CO2 in the inner disk, which is negligible compared to gas phase production (Bosman et al., 2017). Considering the latter, OH can react with CO to form CO2 via: CO + OH CO2 + H (Smith et al., 2004), which has a negligible activation energy and is most effective at gas temperatures up to 400 K, after which destruction channels for OH and CO2 start to become important. In fact, OH can also react with molecular hydrogen to form water: OH + H2 H2O + H (Glassgold et al., 2009). This reaction becomes effective at 200–300 K, due to a high activation barrier of nearly 1700 K, necessary to form OH first (McElroy et al., 2013). Given the exothermic (spontaneous) nature of the CO2 and H2O formation channels, and the availability of gas phase H2 and OH reactants in the warm disk atmosphere (200–1000 K), H2O should be more abundant than CO2 in the inner disk but on condition of neglecting photochemical effects and vertical or turbulent mixing. This was first evident from Spitzer observations, where a low CO2 abundance, 10-4 with respect to H2O ( 10-8 with respect to the total gas density) was found (Salyk et al., 2008; Pontoppidan & Blevins, 2014), and it has been repeatedly confirmed by JWST observations in nearby regions (e.g., Banzatti et al., 2025).
@@
There is, however, mounting evidence of CO2-rich inner disk spectra, at least among T Tauri sources (Gasman et al., 2023; Grant et al., 2023; Vlasblom et al., 2025; Salyk et al., 2025), alongside modeling efforts (Bosman et al., 2022; Sellek et al., 2025) investigating the origin of variations in the spectra of CO2 in contrast to water. XUE 10 is a special case, exhibiting the highest contrast between CO2 and H2O mid-infrared emission to date. It is currently also the only Herbig disk with a CO2-rich mid-infrared spectrum. We therefore need to explain why XUE 10 lacks bright H2O vapor emission, and how CO2 vapor can become so abundant and dominate over H2O at the disk gas temperatures and locations probed at mid-infrared wavelengths. Additionally, we aim to investigate the origin of the potential detection of oxygen isotopic anomalies in CO2 vapor in the inner disk, and contextualize our findings with the evolution of disks around young, intermediate-mass stars (⋆ 1.5 ⊙), and the external environment. We propose the following two explanatory scenarios and discuss their application to the case of XUE 10:
@@
The molecular emission observable in the inner disk is determined by in situ processes, i.e., stellar UV irradiation and local gas phase chemistry.
@@
The observable inner disk emission strongly depends on radial transport processes of gas and dust from the outer disk, whose outcome is not reset by in situ processes. These include radial drift of ice-rich dust grains and gas advection, but also processes that can impede their radial transport, specifically outer disk truncation (e.g., via external UV photoevaporation), and dust traps/gas cavities.
5.2.1 In situ chemistry
@@
The harsher UV irradiation coming from a central Herbig star, compared to a cooler T Tauri star, can drive substantial photo-dissociation of water according to the reaction: H2O + OH + H (Heays et al., 2017). If this reaction proceeds on relatively short timescales compared to vertical or turbulent mixing in the disk and compared to the local gas phase H2O production, water is unable to self-shield and its abundance lowers down, while the freed OH radicals can be sequestrated by CO to form CO2. CO2 itself can undergo photo-dissociation, releasing CO and O via: CO2 + CO + O (Heays et al., 2017), and O2 secondarily (Lu et al., 2014). If the UV photo-dissociation rate of H2O and CO2 molecules were of the same order of magnitude, both species would efficiently self-shield. In this sense, the CO2–H2O dichotomy could be explained by the disk thermal structure. In other words, H2O and CO2 (and OH) would get efficiently destroyed in the upper warm disk layers by stellar irradiation, while in the deeper and cooler layers atomic O and OH molecules would rather form CO2 instead of water (Bosman et al., 2018). However, both H2O and OH (with much fewer detections of CO2) have been observed toward several disks around older A- and B-type Herbig stars in nearby regions (Fedele et al., 2011; Banzatti et al., 2020), where the central stars have comparable luminosities and likely emit stronger UV radiation than the F-type star in XUE 10. We therefore speculate that CO2 photo-dissociation is not as significant as H2O photo-dissociation in XUE 10, and that CO2 but not water may self-shield in its inner regions. This implies that alternatively to thermal effects, a high column density of CO2 can be achieved in the inner disk through gas chemistry by actively removing water. Interestingly, this possibility agrees with a thermochemical DALI synthetic spectrum where water self-shielding is not included (Bosman et al., 2022), pointing toward a deep atmospheric layer of CO2. The gas temperature derived in this work for 12CO2 ( 370 K; see Table 2) together with the observation of its hot -branch bandheads at 13.9 and 16.2 m (excited only in very dense gas), further support the hypothesis of a deep CO2 layer in the disk atmosphere of XUE 10. We note that 12CO2 may be partially non-LTE excited, particularly in the -branch. Non-LTE excitation conditions may arise due to infrared pumping fields, which are known to affect other species such as CO and hydrogen cyanide (HCN) at infrared wavelengths (e.g., Thi et al., 2011; Bruderer et al., 2015). This would imply a lower gas temperature and a larger emitting area than estimated under LTE assumptions. A colder gas reservoir (gas 300 K) would in turn strengthen the hypothesis of an efficient CO2 gas phase production in the inner disk at the expense of water (Bosman et al., 2017; Salyk et al., 2025).
@@
It is important to underline that the production of CO2 vapor through its gas phase formation channel crucially depends on the availability of OH molecules. This radical primarily forms through the reaction O + H2 OH + H, with an activation barrier of 1700 K (McElroy et al., 2013), through water photo-dissociation, and in a minor fraction via H + O2 OH + H, which is characterized by a much higher activation energy of 8000 K (Baulch et al., 1992). The non-detection of OH in XUE 10 further suggests a low water column density in its inner disk, which may limit the photo-dissociation-driven replenishment of this radical (Zannese et al., 2024). The rotationally excited emission lines of OH may also remain undetectable with JWST due to limited sensitivity above the dust continuum.
@@
This in situ scenario can explain a low H2O column density, the non-detection of OH, and the observation of a high column density of CO2, and therefore the detection of bright CO2 isotopologues in emission. However, as water must be removed to allow for efficient CO2 production, this scenario critically depends on the relative timescales of H2O photo-dissociation and of vertical mixing in the inner disk of XUE 10. Concerning the latter, observational evidence suggests weak turbulence in the well-studied Herbig disk HD 163296 (Flaherty et al., 2015), with no significant differences in turbulent mixing strength reported between Herbig and T Tauri disks (Mulders & Dominik, 2012). The oddity might therefore regard the timescale over which H2O photo-dissociates, which must be unusually short in XUE 10 in comparison to Herbig disks of earlier spectral type (A and B), and more evolved ( 3 Myr). Moreover, the presence of potential oxygen isotopic anomalies in CO2 vapor in the inner disk cannot be explained solely by local gas phase channels, unless a substantial dust shielding is present in the inner disk, preventing the UV photo-dissociation of the optically thinner CO2 isotopologues forming in the gas phase. We propose that a more realistic scenario should include the impact of radial transport processes of gas and ice-rich dust grains from the outer to the inner disk, which we consider in the following section.